OpenStudy (anonymous):

1. What is the solution to the equation? 7x + 6 = –22 A.x= -64/7 B.x = –4 C. x= -16/7 D.x = 4

OpenStudy (haichi):

The Answer is (B). The reason is because you must subtract 6 from both sides then divide by 7 on both sides

OpenStudy (haichi):

Since you subtract 6 from both sides the equation is 7x=-28

OpenStudy (anonymous):

kk thx can u help with 4 more ?

OpenStudy (haichi):

Sure :D

OpenStudy (anonymous):

What is the solution to the equation? 12y + 6 = –18 A. y = –2 B. y = –1 C. y = 1 D. y = 2

OpenStudy (haichi):

(A), It is just the same principle as I have explained previously

OpenStudy (haichi):

12y=-24

OpenStudy (anonymous):

What is the solution to the equation? 3-z/4=8 A. z = –44 B. z = –20 C. z = 20 D. z = 44

OpenStudy (haichi):

Z=-20

OpenStudy (anonymous):

What is the solution to the equation? 4 + 3s – 9s = 22 A. s = –3 B. s = –1.5 C. s = 1.5 D. s=4 2/3

OpenStudy (haichi):

A

OpenStudy (anonymous):

What is the solution to the equation? 2q + 18 = –5q – 3 A. q = –5 B. q = –3 C. q = 3 D. q = 5 last oneee

OpenStudy (haichi):

B

OpenStudy (anonymous):

thx so much it got 100 %